OpenStudy (anonymous):

solve for the derivative og the folowing function implicitly x^3 - 3x^2y + 3xy^2 = 12 help me to solve this.

OpenStudy (anonymous):

@triciaal help

OpenStudy (triciaal):

I am testing my memory

OpenStudy (anonymous):

i have example here but i dont know how to do it?

OpenStudy (anonymous):

D(x^3 - xy + y^2) =D(7) 3x^2 - ((1)y + y' x) + 2y y' = 0 3x^2 - (y + y' x) + 2y y' = 0 3x^2 - y - y'x + 2yy' = 0 collecting like terms, - y'x + 2yy'= y- 3x^2 y' (2y-x) = y- 3x^2 y' =(y- 3x^2) /(2y-x)

OpenStudy (triciaal):